OpenStudy (anonymous):

∫e^asin^-1x dx/√ 1-x^2 .. it seems little hard to me to do.

ganeshie8 (ganeshie8):

u sub wil do this time also... start, it wil get easy :)

ganeshie8 (ganeshie8):

maybe il get u started a bit : u = asin^-1x

OpenStudy (anonymous):

let,e^asin^-1x=z or,d/dx e^asin^-1x=dz/dx or,a e^asin^-1x?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

try something like below :- let, asin^-1x=z or,d/dx asin^-1x=dz/dx a/√ 1-x^2 = dz/dx ...

OpenStudy (anonymous):

okay thank you :)

ganeshie8 (ganeshie8):

np :)